# arXiv:2603.02185v1[cond-mat.mtrl-sci] 02 Mar 2026

### v1

@@ 4 Supporting Information
The H4, F’4.5, and H18 configurations defined in this work consist of $3\times 3$, $4\times 2$, and $2\times\sqrt{3}$ MoS2 supercells, respectively. Before plotting the band structures, we check how the Brillouin zones are folded. In Figure 5, we present the corresponding positions of the K(K’) point where the direct gap of a primitive MoS2 monolayer is located.
Figure 5: K-point mapping in (a) H4, (b) F’4.5, and (c) H18 configurations. The first Brillouin zone of primitive MoS2 is illustrated by dashed lines, and the folded ones are shown by solid lines. The red lines illustrate the high-symmetry paths along which the band structures are plotted.
@@
To complete the investigation of the coverage dependent properties, the cases with two and three horizontally adsorbed anthracenes (F2 and F3) in the $6\times 6$ supercell of MoS2 are constructed, as shown in Figure 6.
Figure 6: Top views of the heterostructures with (a) F2 and (b) F3 configurations, created using the interlayer distance of the F1 configuration.
@@
In this work, the $GW_{0}$ quasiparticle corrections are taken at the $G_{1}W_{0}$ level. We check its validity by iterating the Green function $G$. As shown in Figure 7, $G_{1}W_{0}$ is sufficient for the investigation of level alignments. Since this convergence test is done for H18 where the relative position of anthracene VBM has the largest change from $G_{0}W_{0}$ to $G_{1}W_{0}$ among all heterostructures, we conclude that the calculation of level alignments of such hybrid systems requires at least $G_{1}W_{0}$.
Figure 7: The position of anthracene VBM (top of the HOMO band) calculated by partially self-consistent $GW_{0}$: from $G_{0}W_{0}$ to $G_{5}W_{0}$. The MoS2 VBMs of all cases are moved to 0 eV.

### v2

# arXiv:2603.02185v2[cond-mat.mtrl-sci] 03 Mar 2026

@@ 4 Supporting information
The H4, F’4.5, and H18 configurations defined in this work consist of $3\times 3$, $4\times 2$, and $2\times\sqrt{3}$ MoS2 supercells, respectively. Before plotting the band structures, we check how the Brillouin zones are folded. In Figure S1, we present the corresponding positions of the K(K’) point where the direct gap of a primitive MoS2 monolayer is located.
Figure S1: K-point mapping in (a) H4, (b) F’4.5, and (c) H18 configurations. The first Brillouin zone of primitive MoS2 is illustrated by dashed lines, and the folded ones are shown by solid lines. The red lines illustrate the high-symmetry paths along which the band structures are plotted.
@@
To complete the investigation of the coverage dependent properties, the cases with two and three horizontally adsorbed anthracenes (F2 and F3) in the $6\times 6$ supercell of MoS2 are constructed, as shown in Figure S2.
Figure S2: Top views of the heterostructures with (a) F2 and (b) F3 configurations, created using the interlayer distance of the F1 configuration.
@@
In this work, the $GW_{0}$ quasiparticle corrections are taken at the $G_{1}W_{0}$ level. We check its validity by iterating the Green function $G$. As shown in Figure S3, $G_{1}W_{0}$ is sufficient for the investigation of level alignments. Since this convergence test is done for H18 where the relative position of anthracene VBM has the largest change from $G_{0}W_{0}$ to $G_{1}W_{0}$ among all heterostructures, we conclude that the calculation of level alignments of such hybrid systems requires at least $G_{1}W_{0}$.
Figure S3: The position of anthracene VBM (top of the HOMO band) calculated by partially self-consistent $GW_{0}$: from $G_{0}W_{0}$ to $G_{5}W_{0}$. The MoS2 VBMs of all cases are moved to 0 eV.